The centre of regular hexagonis at the point Z=i  if one of its vertices is at 2+i , then the adjacent vertices of  2+i are at the points                                a)1+2i       b)i+1+3½         c)2+i+(1+31/2)                                                                                                    d)1+i(1+31/2)         e)1-i(1+31/2)

Grade:

1 Answers

venkateshwarrao
7 years ago
Given center Z° of the regular hexagon.Z1=2+I is the given vertex,let Z2 and Z6 be the adjacent vertices.Z2=Zo+(Z1-Zo)e^2πi/nZ2= i+(2+i-i)(1/2+i√3/2)=i+1+i√3=1+i(1+√3). similarlyZ6=1-i(1+√3). {Note:(Zn=Zo+(Z1--Zo)e^-(2πi/n}.